Vitex, Inc. manufactures a popular consumer product and it has provided the following data excerpts from its standard cost system: Inputs (1) Standard Quantity or Hours (2) Standard Price or Rate Standard Cost (1) × (2) Direct materials 2.10 pounds $ 16.30 per pound $ 34.23 Direct labor 1.00 hours $ 15.80 per hour $ 15.80 Variable manufacturing overhead 1.00 hours $ 9.30 per hour $ 9.30 Total standard cost per unit $ 59.33 Total Variances Reported Standard Cost* Price or Rate Quantity or Efficiency Direct materials $ 547,680 $ 10,324 F $ 32,600 U Direct labor $ 252,800 $ 3,400 U $ 15,800 U Variable manufacturing overhead $ 148,800 $ 5,000 F $ ?† U *Applied to Work in Process during the period. The company's manufacturing overhead cost is applied to production on the basis of direct labor-hours. All of the materials purchased during the period were used in production. Work in process inventories are insignificant and can be ignored. Required: 1. How many units were produced last period? 2. How many pounds of direct material were purchased and used in production? 3. What was the actual cost per pound of material? (Round your answer to 2 decimal places.) 4. How many actual direct labor-hours were worked during the period? 5. What was the actual rate paid per direct labor-hour? (Round youranswer to 2 decimal places.) 6. How much actual variable manufacturing overhead cost was incurred during the period?Case 10-17 Working Backwards from Variance Data (L010-1, LO10-2, LO10-3] Vitex, Inc. manufactures a popular consumer product
Answer #1

Solution:

1)No of units produced = Total standard cost /standard cost per unit = $547,680 / 34.23 = 16,000

2)Pounds of direct materials purchased and used in production :

Materials quantity variance =(standard quantity -actual quantity ) standard price

-32,600 = (16,000 *2.10 - X)16.30

-32,600 = (33,600 - X)16.30

-32,600 = 547,680 - 16.30X

16.30X = 580,280

X = 35,600

3)Actual cost per pound

Materials price variance =(standard price - actual price)actual quantity

10,324 = (16.30 - X)35,600

10,324 = 580,280 -35,600X

35,600X = 569,956

X = 16.01

4)Calculate actual direct labor hours worked:

Direct labor efficiency variance =(standard hours - actual hours)standard rate

-15,800 =(16,000 - X)15.80

-15,800 = 252,800 - 15.80X

15.80X = 268,600

X=17,000 Hours

5)

Actual rate per direct labour hour:

Labour rate variance =(Standard rate - actual rate)Actual hours

-3,400 =(15.80 - X)17,000

-3,400 = 268,600 - 17,000X

17,000X = 272,000

X=16 per hour

6)

Actual variable overhead cost:

Actual variable overhead = standard overhead + unfavorable variance

=148,800 +4,300

=153,100

Note:

Unfavorable variance = Standard DLH - actual DLH

=16,000 - 17,000

=-1,000

Now multiply 1000 * standard overhead rate 9.30

-1,000 * 9.30 = -9300

Now add -9300 to the favorable rate 5,000

=-9,300 + 5000

=4,300